### v1

@@ Abstract
Time-series with time-varying variance pose a unique challenge to uncertainty quantification (UQ) methods. Time-varying variance, such as volatility clustering as seen in financial time-series, can lead to large mismatch between predicted uncertainty and forecast error. Building on recent advances in neural network UQ literature, we extend and simplify Deep Evidential Regression and Deep Ensembles into a unified framework to deal with UQ under the presence of volatility clustering. We show that a Scale Mixture Distribution is a simpler alternative to the Normal-Inverse-Gamma prior that provides favorable complexity-accuracy trade-off. To illustrate the performance of our proposed approach, we apply it to two sets of financial time-series exhibiting volatility clustering: cryptocurrencies and U.S. equities.
@@ 1 Introduction
Asset returns are known to exhibit irregular bursts of high volatility that cluster in time (termed volatility clustering; Cont,, 2001). This poses a challenge to practitioners during portfolio construction which involves the trade-off of return and risk. To motivate the discussion, consider the following simple thought experiment. Suppose an investor has a model that can perfectly forecast next day’s asset returns and that the investor’s goal is to maximize terminal wealth. Then, on each day, the most rational decision would be to place all of the investor’s wealth into the asset with the highest expected return on the next day. Next, suppose that the investor’s model is a noisy estimator of future asset returns. Then, the investor may choose to diversify across multiple assets. This has led to the development of various models for optimal bet allocation that depend on some measures of risk, such as Kelly criterion (Kelly,, 1956; Byrnes and Barnett,, 2018) and Bayesian-based portfolio optimization (Black and Litterman,, 1991).
@@
Forecast uncertainty has an important role in many applications. Such quantity is easy to obtain for statistical models such as linear regression. However, classical neural networks for regression problems are typically trained using MSE and provide point estimates for the mean prediction conditional on the input without regards for the conditional variance (see Goodfellow et al.,, 2016). As a modeller (in our case, an investor), one is concerned with predictive uncertainty (Gawlikowski et al.,, 2021). This is the total uncertainty around a point estimate. Predictive uncertainty can be decomposed into (Gruber et al.,, 2023): aleatoric uncertainty, and epistemic uncertainty. Aleatoric uncertainty originates from the stochastic relationship between input variable $X$ taking value $x$ and output variable $Y$ (Gruber et al.,, 2023). As long as the conditional distribution of $Y|x$ is not degenerate (i.e., $Y$ cannot be perfectly predicted), there will always be aleatoric uncertainty. Aleatoric uncertainty does not typically depend on sample size. By contrast, epistemic uncertainty is attributable to the model and typically scales inversely with sample size (Meinert et al.,, 2022). Epistemic uncertainty can be further decomposed into model uncertainty, which relates to the correct specification of the model, and parametric uncertainty, which relates to the correct estimation of model parameters (Sullivan,, 2015; Gruber et al.,, 2023). Epistemic uncertainty refers to the part of predictive uncertainty that is reducible through additional information (e.g., more observations and additional variables). In practice, a clear separation between aleatoric and epistemic uncertainties is often impossible. To illustrate, consider the (fair) dice rolling experiment, commonly considered to be a process of pure randomness. However, if the initial position and each rotation of the dice can be measured, then it is possible to predict the outcome of each dice roll (Hora,, 1996; Gruber et al.,, 2023). Thus, what is truly aleatoric (i.e., unpredictability of dice roll) and what is epistemic (i.e., initial position and rotation of the dice are merely missing variables) may be difficult to disentangle from a philosophical perspective.
Traditionally, neural network uncertainty quantification (UQ) requires the use of Bayesian methods or evaluation of the model in unseen data (Meinert et al.,, 2022). A Bayesian neural network (BNN) is a full probabilistic interpretation of neural network, by placing priors on network weights and inducing a distribution over a parametric set of functions (MacKay,, 1992; Neal,, 1996; Gal,, 2016). Modern BNNs can be trained using Markov Chain Monte Carlo (MCMC, e.g., the Metropolis-Hastings algorithm; Hastings,, 1970) and Variational Inference techniques (Jospin et al.,, 2022). Jospin et al., (2022) notes four advantages of using BNNs over classical neural networks, with two being relevant to UQ. First, Bayesian methods provide a natural approach to UQ and are better calibrated than classical neural networks (Mitros and Namee,, 2019; Kristiadi et al.,, 2020; Ovadia et al.,, 2019; Jospin et al.,, 2022). Second, BNN allows distinguishing between epistemic uncertainty and aleatoric uncertainty. However, despite their advantages, MCMC-based methods are computationally expensive (Quiroz et al.,, 2019). Thus, limiting the applicability of BNNs.
Recent advances (see Gawlikowski et al.,, 2021 for a recent survey) have focused on predicting the conditional distribution that is most likely to have generated the data and thus bridging the gap between BNNs and classical neural networks. In particular, using a neural network to generate parameters of a conditional distribution that is assumed to have generated the data (Lakshminarayanan et al.,, 2017; Amini et al.,, 2020) offers an attractive trade-off between adequately quantifying uncertainty and avoiding the computational cost of a full Bayesian treatment. In Lakshminarayanan et al., (2017) (the Ensemble method, also know as Deep Ensembles), regression target $y$ is assumed to be drawn from $y\sim\operatorname{N}(\mu,\sigma^{2})$, where $\operatorname{N}$ is the Normal distribution, $\mu$ is the expectation of $y$ and $\sigma^{2}$ models aleatoric uncertainty. In this setup, $\sigma^{2}$ is incapable of quantifying epistemic uncertainty. Lakshminarayanan et al., (2017) addressed this by using an ensemble of neural networks with randomly initialized weights. Each network settles in a different local minima and produces different $\mu$ and $\sigma^{2}$ for the same input. The variance of $\mu$ across the ensemble thus provides an estimate of epistemic uncertainty. Addressing this shortcoming, Amini et al., (2020) (the Evidential method, also know as Deep Evidential Regression) proposed to place an evidential prior [^1] In contrast to conventional priors in Bayesian inference where the modeller has to specify the parameters of the prior distribution, the evidential prior (e.g., NIG in Evidential) learns these hyperparameters from the data. Note that NIG is a conjugate prior to the Normal distribution (Bernardo and Smith,, 2000)., the Normal-Inverse-Gamma distribution (NIG), on $\mu,\sigma^{2}$. In this construct, prediction $\mu$ is assumed to be drawn from the priors: $\mu\sim\operatorname{N}(\gamma,\sigma^{2}\nu^{-1})$ and $\sigma^{2}\sim\operatorname{InvGam}(\alpha,\beta)$, where $\sigma^{2}$ remains as an estimate of aleatoric uncertainty, $\operatorname{InvGam}$ (or $\operatorname{IG}$) is the Inverse-Gamma distribution and $\sigma^{2}\nu^{-1}$ (with $\sigma^{2}\sim\operatorname{InvGam}(\alpha,\beta)$) is estimated epistemic uncertainty. Epistemic uncertainty is linked to aleatoric uncertainty via $\nu$, which is learnt from the data. The marginal distribution of a Normal likelihood with NIG prior is the Student’s t-distribution. This mimics a Bayesian setup and circumvents the costly computational burden of MCMC methods by analytically integrating out unobserved variables. Ensemble and Evidential require only minimal modifications to a conventional neural network architecture — requiring only the negative log-likelhood (NLL) function of the marginal distribution as loss function and a new output layer. Evidential has been applied to navigation (Liu et al.,, 2021; Cai et al.,, 2021; Singh et al.,, 2022) and medical fields (Soleimany et al.,, 2021; Li and Liu,, 2022), and has been extended into the multi-task learning domain (Oh and Shin,, 2022). Multivariate models related to Evidential include the Natural Posterior Network which also uses a conjugate prior (NIG for regression problems and Dirichlet for categorical classification problems; Charpentier et al.,, 2021), and Regression Prior Networks which uses a Normal-Wishart prior (Malinin et al.,, 2020).
However, more recent works have highlighted weaknesses of the Evidential method. Scoring rules are a class of loss functions that measure the discrepancy between a predicted distribution and the observed distribution (Gneiting and Raftery,, 2007). A scoring rule is proper if the score is maximized when the discrepancy is minimized, and is strictly proper if the maximum is unique. Thus, strictly proper scoring rules provide attractive loss functions for scoring probabilistic forecasts. Evidential can be interpreted as a hierarchical method with a prior distribution that controls the data distribution. Bengs et al., (2023) argues that in order for hierarchical methods such as Evidential to comply with the requirements of proper scoring rules, rather than training on observable values of $y$, the predictor must be trained on the imaginary distribution around each observation that depicts its uncertainty, which cannot possibly exist. This requirement stems from the definition of proper scoring, which requires the learner be scored against the “ground truth”. As $\alpha$ and $\beta$ relate to the prior distribution in Evidential, they are not directly observed as data. Thus, hierarchical methods lack theoretical guarantee on the robustness of their estimated distributions. Similarly, Meinert et al., (2022) argued that unlike aleatoric uncertainty, epistemic uncertainty has no “ground truth” and is difficult to estimate objectively. To motivate this argument, Meinert et al., (2022) used the example of points lined up perfectly in a straight line. If one point is perturbed such that the points no longer form a straight line. Without relying on a-priori assumptions, it is impossible to perform point-wise separation of aleatoric and epistemic uncertainties (i.e., whether the single deviation is due to noise or correctness of the linear model and its estimated slope). The marginal t-distribution of NIG is overparameterized, which leads to the finding that it is possible to minimize the NLL irrespective of $\nu$ (interpreted as “strength of the data” in Amini et al.,, 2020). As a further critique of the network architecture, we note that all four hyperparameters of Evidential are derived from the same latent representation outputted by the last hidden layer. The four hyperparameters can have vastly different scales (e.g., in our motivating application, $\gamma$ is in scale of 0.01, while $\nu$ is in scale of 10). We consider this feature to be a weakness of these approaches as the latent representation has to provide a sufficiently rich encoding to linearly derive all hyperparameters of the distribution. Nonetheless, successful applications of Evidential on real world datasets has led Meinert et al., (2022) to conclude that Evidential is a heuristic to Bayesian methods and may be appropriate for applications that aim to capture both aleatoric and epistemic uncertainties but do not demand an accurate distinction between them, such as our motivating application.
In this work, we are concerned with neural network UQ for time-series that exhibit time-varying variance, such as time-series of asset returns. We combine and extend Ensemble and Evidential into a framework (the Combined method) for quantifying forecast uncertainty of this class of time-series. We propose to formulate the problem using the scale mixture distribution (SMD), a simpler alternative to the NIG prior, to address some of the shortcomings highlighted by Meinert et al., (2022) and Bengs et al., (2023). In SMD, a sole Gamma prior is placed on the scaling factor of variance of the Normal distribution, rather than both the mean and variance in Evidential. This is motivated by our asset return forecasting application, where the mean is typically close to zero (in scale of 0.01) and thus uncertainty is negligible, and volatility is significantly larger (standard deviation in scale of 0.1). This is consistent with fitting a return series with models such as Generalized ARCH (GARCH; Bollerslev,, 1986) in which the mean process is typically assumed zero or first order autoregressive (Carroll and Kearney,, 2009). Integrating out the scaling factor of SMD results in a marginal t-distribution, of which its variance indicates the predictive uncertainty. Epistemic uncertainty is assumed to be the difference between variance of the marginal t-distribution and variance of the assumed Normal data distribution. This simplification trades off granular attribution of aleatoric and epistemic uncertainties afforded by the NIG prior but allows the reduction of the number of effective parameters by one and resolves the overparameterization of NIG, as highlighted by Meinert et al., (2022). We also propose a novel architecture to model parameters of the marginal distribution using disjoint subnetworks, rather than a single output layer as in Ensemble and Evidential. We show through an ablation study in Section 4.2 that this is crucial to forecasting predictive uncertainty that closely tracks forecast error when the time-series exhibit volatility clustering. As both forecast accuracy and estimation of predictive uncertainty are important to our motivating application, we incorporate model averaging into our Combined method and show that it significantly improves forecast accuracy without significantly changing the estimated predictive uncertainty. This work also provides a template for UQ in time-series that exhibit volatility clustering, such as time-series of asset returns.
@@
To illustrate our contributions, we apply our proposed method to cryptocurrency and U.S. equities time-series forecasting. Cryptocurrencies are an emerging class of digital assets. They are highly volatile and frequently exhibit price bubbles (Fry and Cheah,, 2016; Hafner,, 2018; Chen and Hafner,, 2019; Núñez et al.,, 2019; Petukhina et al.,, 2021), with large volumes of high frequency data (e.g., prices in hourly intervals) freely available from major exchanges. This makes cryptocurrencies an ideal testbed for UQ methodologies in financial applications. Given the extreme levels of volatility, we view cryptocurrencies as one of the most challenging datasets for this type of application. A comparison in U.S. equities is also provided which illustrates performance in conventional financial time-series. In the rest of this paper, we first describe the setup of our motivating application (asset return forecasting) in Section 2.1 and review of related works in Section 2.2. We describe our proposed framework in Section 3. Data description and empirical results of applying Ensemble, Evidential and Combined on cryptocurrency and U.S. equities are detailed in Section 4. Whilst this paper is focused on UQ in time-series that exhibit volatility clustering, in Appendix A.1, we also provide a direct comparison to Evidential and Ensemble using the University of California Irvine Machine Learning Repository (UCI) benchmark datasets (non-time-series), as previously analysed in Hernández-Lobato and Adams, (2015), Gal and Ghahramani, (2016), Lakshminarayanan et al., (2017), and Amini et al., (2020). Finally, concluding remarks are provided in Section 5.
2 Preliminaries

### v2

@@ Abstract
Time-series with volatility clustering pose a unique challenge to uncertainty quantification (UQ) for returns forecasts. Methods for UQ such as Deep Evidential regression offer a simple way of quantifying return forecast uncertainty without the costs of a full Bayesian treatment. However, the Normal-Inverse-Gamma (NIG) prior adopted by Deep Evidential regression is prone to miscalibration as the NIG prior is assigned to latent mean and variance parameters in a hierarchical structure. Moreover, it also overparameterizes the marginal data distribution. These limitations may affect the accurate delineation of epistemic (model) and aleatoric (data) uncertainties. We propose a Scale Mixture Distribution as a simpler alternative which can provide favorable complexity-accuracy trade-off and assign separate subnetworks to each model parameter. To illustrate the performance of our proposed method, we apply it to two sets of financial time-series exhibiting volatility clustering: cryptocurrencies and U.S. equities and test the performance in some ablation studies.
@@ 1 Introduction
1.1 Motivation
In finance, volatility estimation is essential for investors, traders, and financial institutions for risk management serving as a crucial metric for assessing and predicting the degree of price fluctuations in assets. Forecast uncertainty is an important determinant of the optimal trade-off between return and risk, and has long been used in portfolio construction, such as in Kelly criterion (Kelly,, 1956; Byrnes and Barnett,, 2018) and Bayesian-based portfolio optimization (Black and Litterman,, 1991).
Recently, neural network emerged as an optimal method for diverse regression problems due to its remarkable capacity for linear and non-linear mapping, surpassing the constraints of rigid model structures in traditional statistical models. In this paper, we are concerned with neural network UQ in financial time-series that exhibit volatility clustering (Cont,, 2001) — irregular bursts of high volatility that cluster in time. This poses a challenge to practitioners wishing to use neural networks to forecast both asset returns and uncertainties, as irregular bursts of volatility can induce forecast errors. Moreover, challenges also stem from modelling these irregular bursts of volatility using classical neural networks which are typically trained using mean squared error (MSE) and provide point estimates for the target variable $Y$ using the mean prediction $\mu$ conditional on the input $X$ taking value $x$. As MSE considers only the difference of the predicted mean $\mu$ from the observed value of $Y$, it is incapable of estimating the conditional variance $\sigma^{2}$ which refers to the uncertainty of the noisy data $Y$ in UQ (Goodfellow et al.,, 2016).
1.2 Uncertainty quantification
This conditional variance $\sigma^{2}$ originates from the stochastic relationship between the input variable $X$ and output variable $Y$ (Gruber et al.,, 2023). It is termed aleatoric uncertainty — uncertainty in the data that is irreducible through additional information (i.e., more observations). As long as the distribution of $Y|x$ is not degenerate (i.e., $Y$ cannot be perfectly predicted using $X$), there will always be aleatoric uncertainty and is indicated by the variance of $Y|x$. It can be estimated by assigning a distribution that is most likely to have generated the data $Y|x$.
@@ 1.2 Uncertainty quantification
In addition to aleatoric uncertainty, there is also uncertainty attributable to the model. This source of uncertainty is known as epistemic uncertainty. Epistemic uncertainty can be further decomposed into model uncertainty, which relates to the correct specification of the model, such as assumption of normality or linear relationship betwen $Y$ and $X$, and parametric uncertainty, which relates to the correct estimation of model parameters such as mean $\mu$ and variance $\sigma^{2}$ (Sullivan,, 2015; Gruber et al.,, 2023). Epistemic uncertainty is reducible through additional information (e.g., more observations and additional variables) and typically scales inversely with sample size (Meinert et al.,, 2022). In practice, a clear separation between aleatoric and epistemic uncertainties is often impossible. To illustrate, consider the (fair) dice rolling experiment, commonly considered to be a process of pure randomness. However, if the initial position and each rotation of the dice can be measured, then it is possible to predict the outcome of each dice roll (Hora,, 1996; Gruber et al.,, 2023). Thus, what is truly aleatoric (i.e., unpredictability of dice roll) and what is epistemic (i.e., initial position and rotation of the dice are merely missing variables) may be difficult to disentangle from a philosophical perspective.
As a modeller (in our case, an investor), one is concerned with predictive uncertainty (Gawlikowski et al.,, 2021) which is the total uncertainty around a point estimate. Predictive uncertainty is comprised of aleatoric uncertainty, and epistemic uncertainty (Gruber et al.,, 2023). Together, these two sources of uncertainty give a complete picture of the uncertainty of the forecasts [^1] We provide a more in-depth discussion of the types of uncertainty in Sections 2.3.1 and 2.4.1.. To account for both sources of uncertainty, a full Bayesian approach not only assumes a data distribution but also assigns prior distributions to all model parameters. In contrast, the classical likelihood approach considers only the data distribution, often leading to criticisms for underestimating the true level of uncertainty. Similarly, the Bayesian neural network (BNN), where prior distributions are placed on network weights, are able to capture both sources of uncertainty. Simpler BNNs can be trained using Markov Chain Monte Carlo (MCMC) which offers accurate estimates of the posterior distribution, while more complex BNNs can be trained using variational inference which trades off accuracy for tractability (Jospin et al.,, 2022). However, a full BNN is often both computationally costly and requires modification of the network architecture.
1.3 Research background
Recent advances (see Gawlikowski et al.,, 2021 for a recent survey) focuses on predicting the conditional distribution of $Y|x$ by estimating the parameters of the distribution, including both mean $\mu$ and variance $\sigma^{2}$ using the negative log-likelihood (NLL) loss function instead of MSE. This type of neural network is called density network (Bishop,, 1994) and provides an attractive alternative to BNN, as it allows priors to be placed on parameters of the assumed data distribution. Evaluation of the marginal distribution NLL is done by integrating out the prior (i.e., uncertainty of the model parameters), in the spirit of posterior sampling using MCMC for BNNs. Hence, this approach offers an attractive trade-off between adequately quantifying uncertainty as well as avoiding the computational cost of a full Bayesian treatment. Prominent methods in this category include Deep Ensemble (Lakshminarayanan et al.,, 2017; the Ensemble method) and Deep Evidential (Amini et al.,, 2020; the Evidential method). Between the two methods, Ensemble assumes a Normal data distribution with mean and variance parameters, $(\mu,\sigma^{2})$, where $\sigma^{2}$ indicates aleatoric uncertainty. Epistemic uncertainty is estimated by using an ensemble of randomly initialized networks, where each network settles at a different local minima and thus giving different mean and variance estimates. Hence, these estimates reveal parameter uncertainty.
By contrast, Evidential improves on this by assuming a Normal prior for the mean parameter $\mu$ and an Inverse-Gamma prior for the variance parameter $\sigma^{2}$ of the normal data distribution. This is known as the Normal-Inverse-Gamma (NIG) prior. The resultant marginal distribution after integrating out the priors of $\mu$ and $\sigma^{2}$ is a Student’s t-distribution parameterized by parameters of the NIG priors (Bernardo and Smith,, 2000). This Student’s t-distribution is a hierarchical distribution consisting of the Normal data distribution conditional on parameters $\mu,\sigma^{2}$, and the prior distribution for each of $\mu$ and $\sigma^{2}$. As t-distribution has heavier tails than Normal distribution, this hierarchical distribution provides a principled way of capturing epistemic uncertainty without the need for ensembling. Moreover, Evidential requires only minimal modifications to the network architecture, specifically a new output layer and training with the NLL of the marginal distribution.
However, Bengs et al., (2023) argued that hierarchical models, such as Evidential, cannot be well calibrated as the prior distributions are assigned to the latent model parameters $\mu,\sigma^{2}$ which cannot be observed directly. Thus, hierarchical models lack theoretical guarantee on the robustness of their estimated distributions. Moreover, Meinert et al., (2022) argued that the marginal Student-t distribution of the Evidential method is overparameterized. A detailed explanation is provided in Section 2.3.2. Hence, epistemic uncertainty is difficult to estimate objectively. In addition to the shortcomings from a statistical modelling perspective, we also note an architectural shortcoming. In both Ensemble and Evidential methods, output of the final hidden layer is fed into a linear output layer, which outputs two or four hyperparameters of the Normal or marginal t-distribution [^2] In the case of Evidential’s marginal t-distribution, softplus is applied to the output for parameters $\nu,\alpha,\beta$., respectively. We consider this a weakness of these approaches as the latent representation has to provide a sufficiently rich encoding to linearly derive all hyperparameters of the distribution.
Recent works have extended Evidential into various related areas. Hüttel et al., (2023) extended Evidential into nonparametric quantile regressions by replacing the Normal data distribution with an Asymmetric Laplace (AL) distribution with NIG prior. By expressing AL into a mean-variance mixture of Normal distribution and an exponentially distributed mixing variable $z_{i}$, they obtained a marginal t-distribution conditional on $z_{i}$ that is similar to the Evidential method. Other works have focused on regions where the neural network is uncertain due to the lack of training observations, called high uncertainty areas. In Evidential, Amini et al., (2020) proposed to regularize the network with a total evidence penalty added to the marginal NLL to encourage the network to output lower evidence ($\nu$, to be introduced in Section 2.3), and thus greater uncertainty. In this high uncertainty region, Ye et al., (2024) argued that the gradient of the marginal NLL of Evidential is close to zero during training when $\alpha\to 1$ (a parameter of the NIG prior). They proposed to add another regularization term in addition to the total evidence penalty to address the zero gradient problem and showed that the proposed regularization achieved comparable performance to Evidential on the University of California Irvine Machine Learning Repository (UCI) benchmark dataset. Similarly focusing on sparse training samples, Deng et al., (2023) proposed to use the Fisher Information Matrix to up-weight uncertain observations for classification problems. Separately, Meinert and Lavin, (2022) proposed the Multivariate Deep Evidential regression, using a Normal-Inverse-Wishart prior as a multivariate extension of NIG prior for multivariate regressions. Importantly, Meinert and Lavin, (2022) argued that the total evidence penalty is only necessary as the NIG prior is overparameterized. As a workaround, Meinert and Lavin, (2022) suggested to fix $\nu=c\beta$ to reduce the number of parameters of the NIG prior from four to three, where $c$ is a hyperparameter of the model [^3] Meinert and Lavin, (2022) used notations $\nu=r\kappa$, where $\kappa$ maps to $\beta$ in our notations.. While these advances extend the statistical model framework and address some shortcomings of Evidential, we propose an alternative formulation with a simplified statistical model structure which alleviates the need for a total evidence penalty, while also offering potential research directions to nonparametric quantile regression and multivariate regression.
1.4 Proposed method
To tackle the shortcomings highlighted by Meinert et al., (2022), Bengs et al., (2023) and others, we combine and extend the ideas of Ensemble and Evidential methods into a framework called the Combined method for quantifying forecast uncertainty of financial time-series. Specifically, we propose to formulate the data distribution using the scale mixture distribution (SMD), a simpler alternative to the NIG prior, by avoiding prior distributions for the latent mean and variance parameters of the Normal distribution. Specifically, the epistemic uncertainty in SMD is captured by allowing each normally distributed data point to have its own variance $\sigma^{2}\nu^{-1}$, where $\nu$ is the scaling factor. This scaling factor provides flexibility to each data point by enlarging or diminishing the variance for outliers or modal points. Then, a sole Gamma prior is assigned to the scaling factor. Integrating out the scaling factor of SMD also results in a marginal t-distribution and its variance indicates predictive uncertainty, which stems from two sources of uncertainty — the aleatoric uncertainty from the $\sigma^{2}$ of the conditional Normal data distribution and epistemic uncertainty from the Gamma distribution of scaling factor. This simplification reduces the number of effective parameters by one and resolves the overparameterization of NIG, as highlighted by Meinert et al., (2022). This also alleviates the need to regularize the marginal NLL, as noted by Meinert and Lavin, (2022). A detailed explanation is provided in Section 2.4.2.
In addition to statistical model structure, we also propose a novel architecture to model parameters of the marginal distribution using disjoint subnetworks, rather than a single output layer as in Ensemble and Evidential. We show through an ablation study in Section 3.2 that the flexibility offered by the subnetworks is crucial to forecasting predictive uncertainty that closely tracks volatility when the time-series exhibit volatility clustering. As return forecast accuracy is also important in finance, we incorporate model averaging (in the spirit of ensemble) into our Combined method and show that it significantly improves return forecast accuracy without significantly increasing the estimated predictive uncertainty.
@@ 1.4 Proposed method
To illustrate our contributions, we apply our proposed method to cryptocurrency and U.S. equities time-series forecasting. Cryptocurrencies are an emerging class of digital assets. They are highly volatile and frequently exhibit price bubbles (Fry and Cheah,, 2016; Hafner,, 2018; Chen and Hafner,, 2019; Phillip et al.,, 2018; Núñez et al.,, 2019; Petukhina et al.,, 2021), with large volumes of high frequency data (e.g., prices in hourly intervals) freely available from major exchanges. This makes cryptocurrencies an ideal testbed for UQ methodologies in financial applications. Given the extreme levels of volatility, we view cryptocurrencies as one of the most challenging datasets for this type of application. A comparison in U.S. equities is also provided which illustrates performance in conventional financial time-series.
1.5 Contributions and layout
A summary of our contributions is as follows:
  • We propose to place a prior solely on the scaling factor $\nu$ using a SMD, rather than on both mean and variance using the NIG prior. In Table 6 (Appendix B), we compare SMD to NIG using the same network architecture on the UCI dataset. We observe that both RMSE and NLL are overwhelmingly in SMD’s favor, achieving superior results in 6 (of 9) and 8 (of 9) of the datasets for RMSE and NLL, respectively. We attribute this to the simpler formulation of SMD, which resolves the latent variables issue of hierarchical models noted by Bengs et al., (2023) and the overparameterization problem of the NIG prior, noted by Meinert et al., (2022).
  • We show that modelling hyperparameters of the posterior distribution using subnetworks rather than a single output layer (as per Ensemble and Evidential) results in superior performance. We show in the ablation study (Section 3.2, the “Single Output” model in Table 3 and Figure 5) that separate modelling of the hyperparameters is essential in producing predictive uncertainty estimates that closely track actual realized forecast error of Bitcoin and Chevron. We attribute this to the ease of specialization when each hyperparameter of the posterior distribution is modelled through disjoint subnetworks.
  • In additional to the subnetwork modelling of distribution hyperparameters, we show that including squared returns as an additional feature to the long short-term memory (LSTM; Hochreiter and Schmidhuber,, 1997) layers in architectural design is also essential to successfully quantifying uncertainty in financial time-series. We provide a precedence on UQ in financial time-series, a neglected application in UQ literature.
In the rest of this paper, we first describe the setup of our motivating application (asset return forecasting) in Section 2.1 and review related works in Sections 2.2 and 2.3. We describe and discuss our proposed framework in Section 2.4. Data description and empirical results of applying Ensemble, Evidential and Combined methods on cryptocurrency and U.S. equities are detailed in Section 3. Whilst this paper is focused on UQ in time-series that exhibit volatility clustering, in Appendix B, we also provide a direct comparison to Evidential and Ensemble methods using the UCI benchmark datasets (non-time-series), as previously analyzed in Hernández-Lobato and Adams, (2015), Gal and Ghahramani, (2016), Lakshminarayanan et al., (2017), and Amini et al., (2020). Finally, concluding remarks are provided in Section 4.
2 Methodological development in uncertainty quantification